40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Dec 7 11:25:31 2020, gautam, Update, SUS, MC1 suspension glitchy again 
    Reply  Wed Dec 9 16:22:57 2020, gautam, Update, SUS, Yet another round of Sat. Box. switcharoo WFS2.pngWFS_lineNoise.pngWFSchar.pdf
       Reply  Sun Jan 3 16:26:06 2021, Koji, Update, SUS, IMC WFS check (Yet another round of Sat. Box. switcharoo) Screen_Shot_2021-01-03_at_17.14.57.png
    Reply  Thu May 13 11:55:04 2021, Anchal, Paco, Update, SUS, MC1 suspension misbehaving Screenshot_from_2021-05-13_09-50-24.pngMC1_Glitches_Invest2.pdf
       Reply  Thu May 13 19:38:54 2021, Anchal, Update, SUS, MC1 Satellite Amplifier Debugged MC1_UL_Channel_Fixed.png
          Reply  Mon May 24 19:14:15 2021, Anchal, Paco, Summary, SUS, MC1 Free Swing Test set to trigger 
             Reply  Tue May 25 10:22:16 2021, Anchal, Paco, Summary, SUS, MC1 new input matrix calculated and uploaded SUS_Input_Matrix_Diagonalization.pdf
          Reply  Thu Jun 17 11:45:42 2021, Anchal, Paco, Update, SUS, MC1 Gave trouble again SummaryScreenShot.pngMC1_LL_SENSOR_DEAD.png
             Reply  Thu Jun 17 16:37:23 2021, Anchal, Paco, Update, SUS, c1susaux computer rebooted 
                Reply  Tue Jun 22 11:56:16 2021, Anchal, Paco, Update, SUS, ADC/Slow channels issues CDS_FE_Status.png
                   Reply  Tue Jun 22 16:52:28 2021, Paco, Update, SUS, ADC/Slow channels issues shake_and_damp.png
                      Reply  Wed Jun 23 09:05:02 2021, Anchal, Update, SUS, MC lock acquired back again 
                         Reply  Thu Jun 24 16:40:37 2021, Koji, Update, SUS, MC lock acquired back again P_20210624_163641_1.jpg
Message ID: 16219     Entry time: Tue Jun 22 16:52:28 2021     In reply to: 16218     Reply to this: 16222
Author: Paco 
Type: Update 
Category: SUS 
Subject: ADC/Slow channels issues 

After sliding the alignment bias around and browsing through elog while searching for "stuck" we concluded the ITMX osems needed to be freed. To do this, the procedure is to slide the alignment bias back and forth ("shaking") and then as the OSEMs start to vary, enable the damping. We did just this, and then restored the alignment bias sliders slowly into their original positions. Attachment 1 shows the ITMX OSEM sensor input monitors throughout this procedure.


At the end, since MC has trouble catching lock after opening PSL shutter, I tried running burt restore the ioo to 2021/Jun/17/06:19/c1iooepics.snap but the problem persists

Attachment 1: shake_and_damp.png  54 kB  Uploaded Tue Jun 22 17:58:45 2021  | Hide | Hide all
shake_and_damp.png
ELOG V3.1.3-